The primary side of a transformer has 935 turns and is connected to 118 Volts A/C.

The secondary side has 254 turns.

What is the secondary voltage?
Answer #1

N1/N2 = V1/V2

935/254 = 118/V2

V2 = 32 V

Comment in case any doubt please rate my answer....